# Big Varchenko–Gelfand rings and Orbit Harmonics

### v2

@@ 1.1. Chambers and topes
Dorpalen-Barry, Proudfoot, and Wang studied [7] the algebra $\mathrm{VG}$$_{\mathcal{M}}$ of functions on the set ${\mathcal{T}}(\mathcal{M})$ of topes of $\mathcal{M}$. As in the case of arrangements, the algebra $\mathrm{VG}$$_{\mathcal{M}}$ admits a natural filtration via Heaviside functions which gives rise to an associated graded ring $\mathcal{VG}$$_{\mathcal{M}}$. Dorpalen-Barry et. al. established [7, Thm. 1.6] a presentation of $\mathrm{VG}$$_{\mathcal{M}}$ and $\mathcal{VG}$$_{\mathcal{M}}$ as polynomial ring quotients. A topological interpretation of these rings is also proven [7, Thm. 1.1] when $\mathcal{M}$ is as in Example 1.1.
@@ 1.4. Organization
The rest of the paper is organized as follows. Section 2 recalls various notions from (conditional, oriented) matroid theory and makes some easy combinatorial observations which will be used later on. The short Section 3 recasts the work of Dorpalen-Barry, Proudfoot, and Wang via the orbit harmonics of the small locus $\mathcal{Y}_{\mathcal{M}}$. Section 4 is the heart of the paper and contains our main results on the big locus $\mathcal{Z}_{\mathcal{M}}$ and the big graded VG ring $\widehat{{\mathcal{VG}}}_{\mathcal{M}}$. Section 5 describes our results in the special case of the type A braid arrangement and poses a topological open question (Problem 5.1).
@@ Acknowledgements
The author is very grateful to Vic Reiner for ideas which initiated this project and inspiring conversations throughout. The author thanks Ethan Partida and Colin Crowley for pointing out the relationship with matroid Schubert varieties of Remark 4.6. The author was partially supported by NSF Grant DMS-2246846.
@@ 2.3. Restriction and contraction
of signed subsets of $\mathcal{I}-F$. It follows from [2, Lem. 1] that both $\mathcal{M}\mid_{F}$ and $\mathcal{M}^{F}$ are conditional oriented matroids.
@@ 2.4. Circuits and NBC sets
Let ${\mathcal{N}}(\mathcal{M})$ be the family of NBC sets. We have [7, Prop. 4.2] the important coincidence
@@ 3.2. VG rings and orbit harmonics
where $e_{0}(z_{1},\dots,z_{n}):=1$. In light of Proposition 3.2, the following result follows from [7, Thm. 1.6].
@@
It follows from results in [7, Sec. 5] that the set $\{\prod_{i\in N}y_{i}^{+}\,:\,N\in{\mathcal{N}}(\mathcal{M})\}$ of monomials in the $y_{i}^{+}$ indexed by NBC sets descends to a vector space basis of either $S/{\mathbf{I}}(\mathcal{Y}_{\mathcal{M}})$ or $S/{\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Y}_{\mathcal{M}})={\mathbf{R}}(\mathcal{Y}_{\mathcal{M}})\cong{\mathcal{VG}}_{\mathcal{M}}$. Our ‘big’ VG quotients will admit a basis of a similar flavor.
@@ 4.4. Middle quotients for the big locus
Ideals similar to $\widetilde{I}_{\mathcal{M}}$ arise in geometry. If $L\subseteq{\mathbb{C}}^{n}$ is a linear subspace, the matroid Schubert variety is the Zariski closure $Y_{L}\subseteq({\mathbb{P}}^{1})^{n}$ of the image of
@@
Intersecting the coordinate hyperplanes of ${\mathbb{C}}^{n}$ with $L$ gives rise to an oriented matroid $\mathcal{M}$ on the ground set $\mathcal{I}=[n]$. Let $z_{i}\in H^{2}(Y_{L})$ be the Chern class coming from the $i^{th}$ coordinate hyperplane. It follows from work of Huh and Wang [11] that the cohomology ring $H^{*}(Y_{L})$ is generated by $z_{1},\dots,z_{n}$ with relations as in Definition 4.5 together with $z_{i}^{2}=0$ for $i=1,\dots,n$.
@@
$\widetilde{I}_{\mathcal{M}}=(z_{1}\cdot z_{2}\cdot z_{3},\quad z_{4},\quad z_{1}\cdot z_{2}-z_{1}\cdot z_{3},\quad z_{1}\cdot z_{2}-z_{2}\cdot z_{3},\quad z_{1}\cdot z_{3}-z_{2}\cdot z_{3}).$
@@
We check that the two kinds of generators in Definition 4.5 lie in ${\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Z}_{\mathcal{M}})$.
@@
Lemma 2.4 (1) guarantees that the polynomial $\prod_{b\in{\mathcal{B}}}z_{b}-\prod_{b\in{\mathcal{B}}^{\prime}}z_{b^{\prime}}\in\hat{S}$ is homogeneous, and so lies in ${\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Z}_{\mathcal{M}})$. ∎
@@ 4.5. The big orbit harmonics quotient
(Sketch) We start with the case $A={\mathbb{Z}}$. The argument proving Theorem 4.14 (with $\mathcal{L}(\mathcal{M})$-filtrations defined over $\mathbb{Z}$} in the obvious way) goes through to show that the set $\widehat{{\mathcal{N}}}$ spans $\widehat{S}^{\mathbb{Z}}/\widehat{I}^{\mathcal{M}}_{\mathbb{Z}}$ over $\mathbb{Z}$}; the key point is that the relations (4.29)-(4.33) appearing in the proof of Theorem 4.14 involve only the coefficients $\pm 1$. Since any $\mathbb{Z}$}-linear relation on $\widehat{{\mathcal{N}}}$ modulo $\widehat{I}^{\mathbb{Z}}_{\mathcal{M}}$ would induce a $\mathbb{Q}$}-linear relation on $\widehat{{\mathcal{N}}}$ modulo $\widehat{I}^{\mathbb{Q}}_{\mathcal{M}}$, we see that $\widehat{S}^{\mathbb{Z}}/\widehat{I}^{\mathbb{Z}}_{\mathcal{M}}$ is a free $\mathbb{Z}$}-module with $\mathbb{Z}$}-basis $\widehat{{\mathcal{N}}}$. Applying the functor $A\otimes_{\mathbb{Z}}(-)$ gives the proposition. ∎
@@ 5.1. Hilbert series
(5.3) ${\mathrm{Hilb}}({\mathbf{R}}(\mathcal{Z}_{\mathcal{M}_{n}});q)={\mathrm{Hilb}}(\widehat{{\mathcal{VG}}}_{\mathcal{M}_{n}};q)=\sum_{\pi\in\Pi_{n}}q^{n-\#\pi}q\cdot(q+1)\cdots(q+\#\pi-1)$
@@
$n$ ${\mathrm{Hilb}}({\mathbf{R}}(\mathcal{Z}_{\mathcal{M}_{n}};q)={\mathrm{Hilb}}(\widehat{{\mathcal{VG}}}_{\mathcal{M}_{n}};q)$
@@ 5 $150q^{4}+250q^{3}+120q^{2}+20q+1$
We may regard (5.3) as a kind of ‘Stirling distribution’ on the family of ordered set partitions of $[n]$. It may be interesting to study ordered set partition statistics which have this distribution.
@@ 5.2. ${\mathfrak{S}}_{n}$-module structure
as graded ${\mathfrak{S}}_{n}$-modules. The representations (5.4) give a graded deformation of the regular representation of ${\mathfrak{S}}_{n}$. Brauner proved [3, Thm. 4.10] that the graded pieces of (5.4) carry the Eulerian representations of ${\mathfrak{S}}_{n}$ (see also [1]). Brauner’s result holds in the greater generality of coincidental real reflection groups.
@@ 5.3. Topological interpretation?
The cohomology of a space $\widehat{{\mathcal{C}}}_{n}$ solving Problem 5.1 would be concentrated in even degrees with sum of Betti numbers given by the number of ordered set partitions of $[n]$. When $n=2$, the hypothetical space $\widehat{{\mathcal{C}}}_{2}$ would have cohomology presentation
@@
In light of the work of Proudfoot [18] and Moseley [14], Problem 5.1 may be best solved with a space $\widehat{{\mathcal{C}}}_{n}$ equipped with an action of the 1-dimensional torus $T={\mathbb{C}}^{*}$. Let $\widehat{\mathcal{Z}}_{\mathcal{M}_{n}}\subseteq{\mathbb{F}}^{{[n]\choose 2}\times\{+,-,0\}}\times{\mathbb{F}}$ be the variety obtained by placing the locus $\mathcal{Z}_{\mathcal{M}_{n}}$ at height 1 and joining every point in this locus to the origin with a line. The work of [18, 14, 7, 5] inspires hope that the $T$-equivariant cohomology of $\widehat{{\mathcal{C}}}_{n}$ admits presentation $H^{*}_{T}(\widehat{{\mathcal{C}}}_{n})=(\widehat{S}\otimes_{\mathbb{F}}{\mathbb{F}}[t])/{\mathbf{I}}(\widehat{\mathcal{Z}}_{\mathcal{M}_{n}}).$ The work of Dorpalen-Barry, Proudfoot, and Wang [7, Thm. 1.1] motivates an extension of Problem 5.1 to arbitrary real arrangements $\mathcal{A}$} in $V$ and arbitrary open convex sets ${\mathcal{K}}\subseteq V.$

### v3

@@ 1.1. Chambers and topes
Dorpalen-Barry, Proudfoot, and Wang studied [8] the algebra $\mathrm{VG}$$_{\mathcal{M}}$ of functions on the set ${\mathcal{T}}(\mathcal{M})$ of topes of $\mathcal{M}$. As in the case of arrangements, the algebra $\mathrm{VG}$$_{\mathcal{M}}$ admits a natural filtration via Heaviside functions which gives rise to an associated graded ring $\mathcal{VG}$$_{\mathcal{M}}$. Dorpalen-Barry et. al. established [8, Thm. 1.6] a presentation of $\mathrm{VG}$$_{\mathcal{M}}$ and $\mathcal{VG}$$_{\mathcal{M}}$ as polynomial ring quotients. A topological interpretation of these rings is also proven [8, Thm. 1.1] when $\mathcal{M}$ is as in Example 1.1.
@@ 1.4. Organization
The rest of the paper is organized as follows. Section 2 recalls various notions from (conditional, oriented) matroid theory and makes some easy combinatorial observations which will be used later on. The short Section 3 recasts the work of Dorpalen-Barry, Proudfoot, and Wang via the orbit harmonics of the small locus $\mathcal{Y}_{\mathcal{M}}$. Section 4 is the heart of the paper and contains our main results on the big locus $\mathcal{Z}_{\mathcal{M}}$ and the big graded VG ring $\widehat{{\mathcal{VG}}}_{\mathcal{M}}$. Section 5 describes our results in the special case of the type A braid arrangement and poses a topological open question (Problem 5.3).
@@ Acknowledgements
The author is very grateful to Vic Reiner for ideas which initiated this project and inspiring conversations throughout. The author thanks Ethan Partida and Colin Crowley for pointing out the relationship with matroid Schubert varieties of Remark 4.6. The author thanks Jaeseong Oh for pointing out the connection between Equation (5.3) and necklaces. The author was partially supported by NSF Grant DMS-2246846.
@@ 2.3. Restriction and contraction
of signed subsets of $\mathcal{I}-F$. It follows from [3, Lem. 1] that both $\mathcal{M}\mid_{F}$ and $\mathcal{M}^{F}$ are conditional oriented matroids.
@@ 2.4. Circuits and NBC sets
Let ${\mathcal{N}}(\mathcal{M})$ be the family of NBC sets. We have [8, Prop. 4.2] the important coincidence
@@ 3.2. VG rings and orbit harmonics
where $e_{0}(z_{1},\dots,z_{n}):=1$. In light of Proposition 3.2, the following result follows from [8, Thm. 1.6].
@@
It follows from results in [8, Sec. 5] that the set $\{\prod_{i\in N}y_{i}^{+}\,:\,N\in{\mathcal{N}}(\mathcal{M})\}$ of monomials in the $y_{i}^{+}$ indexed by NBC sets descends to a vector space basis of either $S/{\mathbf{I}}(\mathcal{Y}_{\mathcal{M}})$ or $S/{\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Y}_{\mathcal{M}})={\mathbf{R}}(\mathcal{Y}_{\mathcal{M}})\cong{\mathcal{VG}}_{\mathcal{M}}$. Our ‘big’ VG quotients will admit a basis of a similar flavor.
@@ 4.4. Middle quotients for the big locus
The ideals $\widetilde{I}_{\mathcal{M}}$ arise in geometry. Let $L\subseteq{\mathbb{C}}^{n}$ be a linear subspace. Ardila and Boocher introduced [2] the matroid Schubert variety as the Zariski closure $Y_{L}\subseteq({\mathbb{P}}^{1})^{n}$ of the image of
@@
Intersecting the coordinate hyperplanes of ${\mathbb{C}}^{n}$ with $L$ gives rise to a matroid $\mathcal{M}$ on the ground set $\mathcal{I}=[n]$. Let $z_{i}\in H^{2}(Y_{L})$ be the first Chern class of the tautological line bundle over the $i^{th}$ factor in $({\mathbb{P}}^{1})^{n}$, pulled back along $Y_{L}\hookrightarrow({\mathbb{P}}^{1})^{n}.$ Huh and Wang proved [12] that the cohomology ring $H^{*}(Y_{L})$ is generated by $z_{1},\dots,z_{n}$ with relations as in Definition 4.5. We may therefore regard the quotient $\widehat{S}/\widetilde{I}_{\mathcal{M}}$ as the cohomology of $Y_{L}$ with coefficients in ${\mathbb{F}}[y_{i}^{\pm}\,:\,i\in\mathcal{I}]$.
@@
$\widetilde{I}_{\mathcal{M}}=(z_{1}\cdot z_{2}\cdot z_{3},\,z_{4},\,z_{1}\cdot z_{2}-z_{1}\cdot z_{3},\,z_{1}\cdot z_{2}-z_{2}\cdot z_{3},\,z_{1}\cdot z_{3}-z_{2}\cdot z_{3},\,z_{1}^{2},\,z_{2}^{2},\,z_{3}^{2},\,z_{4}^{2}).$
@@
We check that the three kinds of generators in Definition 4.5 lie in ${\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Z}_{\mathcal{M}})$.
@@
Lemma 2.4 (1) guarantees that the polynomial $\prod_{b\in{\mathcal{B}}}z_{b}-\prod_{b\in{\mathcal{B}}^{\prime}}z_{b^{\prime}}\in\hat{S}$ is homogeneous, and so lies in ${\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Z}_{\mathcal{M}})$.
(3) For any $i\in\mathcal{I}$ we have $z_{i}(z_{i}-1)\in{\mathbf{I}}(\mathcal{Z}_{\mathcal{M}})$ so that $z_{i}^{2}\in{\mathrm{gr}}\,{\mathbf{I}}(\mathcal{Z}_{\mathcal{M}})$. ∎
@@ 4.5. The big orbit harmonics quotient
(Sketch) We start with the case $A={\mathbb{Z}}$. The argument proving Theorem 4.14 (with $\mathcal{L}(\mathcal{M})$-filtrations defined over $\mathbb{Z}$}) goes through to show that the set $\widehat{{\mathcal{N}}}$ spans $\widehat{S}^{\mathbb{Z}}/\widehat{I}^{\mathcal{M}}_{\mathbb{Z}}$ over $\mathbb{Z}$}; the key point is that the relations (4.29)-(4.33) appearing in the proof of Theorem 4.14 involve only the coefficients $\pm 1$. Since any $\mathbb{Z}$}-linear relation on $\widehat{{\mathcal{N}}}$ modulo $\widehat{I}^{\mathbb{Z}}_{\mathcal{M}}$ would induce a $\mathbb{Q}$}-linear relation on $\widehat{{\mathcal{N}}}$ modulo $\widehat{I}^{\mathbb{Q}}_{\mathcal{M}}$, we see that $\widehat{S}^{\mathbb{Z}}/\widehat{I}^{\mathbb{Z}}_{\mathcal{M}}$ is a free $\mathbb{Z}$}-module with $\mathbb{Z}$}-basis $\widehat{{\mathcal{N}}}$. Applying the functor $A\otimes_{\mathbb{Z}}(-)$ gives the proposition. ∎
@@ 5.1. Hilbert series
(5.3) ${\mathrm{Hilb}}({\mathbf{R}}(\mathcal{Z}_{\mathcal{M}_{n}});q)={\mathrm{Hilb}}(\widehat{{\mathcal{VG}}}_{\mathcal{M}_{n}};q)=\sum_{\pi\in\Pi_{n}}q^{n-\#\pi}\cdot(q+1)\cdot(2q+1)\cdots((\#\pi-1)q+1)$
@@
$n$ ${\mathrm{Hilb}}({\mathbf{R}}(\mathcal{Z}_{\mathcal{M}_{n}});q)={\mathrm{Hilb}}(\widehat{{\mathcal{VG}}}_{\mathcal{M}_{n}};q)$
@@ 5 $150q^{4}+250q^{3}+120q^{2}+20q+1$
Equation (5.3) has the following combinatorial interpretation.
Proposition 5.1.
For any $n$ we have
(5.4) $\sum_{\pi\in\Pi_{n}}q^{n-\#\pi}\cdot(q+1)\cdot(2q+1)\cdots((\#\pi-1)q+1)=\sum_{k=0}^{n}a_{n,k}\cdot q^{k}$
where $a_{n,k}$ counts unordered (nk)(n-k)-tuples of necklaces with beads labeled by subsets $S\subseteq[n]$ such that the bead labels form a set partition of $[n]$.
Proof.
Recall that we have the generating function identity
(5.5) $\sum_{w\in{\mathfrak{S}}_{n}}q^{n-\text{cyc}(w)}=(q+1)(2q+1)\cdots((n-1)q+1).$
Given a set partition $\pi\in\Pi_{n}$, let ${\mathfrak{S}}_{\pi}\cong{\mathfrak{S}}_{\#\pi}$ be the group of permutations of the blocks of $\pi$. For $w\in{\mathfrak{S}}_{\pi}$, write $\text{cyc}(w)$ for the number of cycles of $w$. For example, if $n=8$ and
$\pi=\{17\mid 25\mid 3\mid 4\mid 68\}$
then one possible element $w\in{\mathfrak{S}}_{\pi}$ is given in cycle notation as
w=(17, 68, 4)(25, 3).w=(17,\,68,\,4)(25,\,3).
In this case we have $\text{cyc}(w)=2$. Restating the identity (5.5) for ${\mathfrak{S}}_{\pi}$ gives
(5.6) $\sum_{w\in{\mathfrak{S}}_{\pi}}q^{\#\pi-\text{cyc}(w)}=(q+1)(2q+1)\cdots((\#\pi-1)q+1).$
Applying (5.6), the summand of Equation (5.4) indexed by $\pi$ may be expressed as
(5.7) $q^{n-\#\pi}\cdot(q+1)\cdot(2q+1)\cdots((\#\pi-1)q+1)=\sum_{w\in{\mathfrak{S}}_{\pi}}q^{n-\text{cyc}(w)}.$
Summing over all $\pi\in\Pi_{n}$ gives the proposition. ∎
Example 5.2.
Let $n=3$. The following family of objects witnesses $a_{3,2}=6$.
123132123132231123
The following family of objects witnesses $a_{3,1}=6$.
$123132231$
Finally, the equality $a_{3,0}=1$ is witnessed as follows.
123
@@ 5.2. ${\mathfrak{S}}_{n}$-module structure
as graded ${\mathfrak{S}}_{n}$-modules. The representations (5.8) give a graded deformation of the regular representation of ${\mathfrak{S}}_{n}$. Brauner proved [4, Thm. 4.10] that the graded pieces of (5.8) carry the Eulerian representations of ${\mathfrak{S}}_{n}$ (see also [1]). Brauner’s result holds in the greater generality of coincidental real reflection groups.
@@ 5.3. Topological interpretation?
The cohomology of a space $\widehat{{\mathcal{C}}}_{n}$ solving Problem 5.3 would be concentrated in even degrees with sum of Betti numbers given by the number of ordered set partitions of $[n]$. When $n=2$, the hypothetical space $\widehat{{\mathcal{C}}}_{2}$ would have cohomology presentation
@@
In light of the work of Proudfoot [19] and Moseley [15], Problem 5.3 may be best solved with a space $\widehat{{\mathcal{C}}}_{n}$ equipped with an action of the 1-dimensional torus $T={\mathbb{C}}^{*}$. Let $\widehat{\mathcal{Z}}_{\mathcal{M}_{n}}\subseteq{\mathbb{F}}^{{[n]\choose 2}\times\{+,-,0\}}\times{\mathbb{F}}$ be the variety obtained by placing the locus $\mathcal{Z}_{\mathcal{M}_{n}}$ at height 1 and joining every point in this locus to the origin with a line. The work of [19, 15, 8, 6] inspires hope that the $T$-equivariant cohomology of $\widehat{{\mathcal{C}}}_{n}$ admits presentation $H^{*}_{T}(\widehat{{\mathcal{C}}}_{n})=(\widehat{S}\otimes_{\mathbb{F}}{\mathbb{F}}[t])/{\mathbf{I}}(\widehat{\mathcal{Z}}_{\mathcal{M}_{n}}).$ The work of Dorpalen-Barry, Proudfoot, and Wang [8, Thm. 1.1] motivates an extension of Problem 5.3 to arbitrary real arrangements $\mathcal{A}$} in $V$ and arbitrary open convex sets ${\mathcal{K}}\subseteq V.$ Remark 4.6 suggests that a solution to Problem 5.3 might involve an enlargement of matroid Schubert varieties.